STORM2 BX-C 01/23/13


  • Back-optics aligned
  • Focus lock working
  • Laser power:
    • 647 laser at 200 mW source: 130 mW at power meter
    • 405 laser: max power = 1.89 mW.  10% = .09 mW
  • z-stage drift without focus lock is ~800 nm per min
    • drift about 900 nm down during the time it takes to record a z-calibration movie.  Utterly unusable.

STORM imaging

  • imaged several sections individually.  120,000 frames without needing 405.  May need to discard substantial number of frames due to too high emitter density.
  • generated overview mosaic on Steve to pick good cell regions
  • focus lock froze during O/N imaging.  — Consequently z-focus was lost and many sections are blank. :(.


  • repeat z-calibration. Prior to calibration, focus holding pretty steady without lock.
This entry was posted in General STORM and tagged . Bookmark the permalink.